# ARMAr-LASSO: Mitigating the Impact of Predictor Serial Correlation on the LASSO

### v2

@@ Appendix C Monte Carlo Experiments
In this Section, we conduct Monte Carlo experiments to assess numerically the approximation of the density of $\widehat{c}_{ij}^{x}$ to $\mathcal{D}(r)$, as described in Section 2.1 of the main text. In particular, we compare the density of $\widehat{c}_{ij}^{x}$ obtained by simulations (indicated as $d(r)$) with both the asymptotic distribution provided in Theorem 2 of the main text (indicated as $\mathcal{D}(r)$) and its estimate, namely, where the parameters are estimated from the sample of 5000 Monte Carlo replications (indicated as $\widehat{\mathcal{D}}(r)$). After, we expand the theoretical results in more generic contexts, relaxing the assumption that the predictors are orthogonal Gaussian AR(1) processes.
C.1 Numerical Approximation of $d(r)$ to $\mathcal{D}(r)$
@@
We generate data from the bivariate process $\mathbf{x}_{t}=\mathbf{D}_{\phi}\mathbf{x}_{t-1}+\mathbf{u}_{t}$ for $t=1,\dots,T$, where $\mathbf{D}_{\phi}$ is a $2\times 2$ diagonal matrix with the same autocorrelation coefficient $\phi$ in both positions along the diagonal, and $u_{t}\sim N(\boldsymbol{0}_{2},\boldsymbol{I}_{2})$. We consider $T=50,100,250$ and $\phi=0.3,0.6,0.9,0.95$ – thus, the parameter $\ddot{\phi}$ in $\mathcal{D}(r)$, here equal to $\phi^{2}$, takes values 0.09, 0.36, 0.81, 0.90. The first row of Figure S.4 (Plots (a), (b), (c)) shows, for various values of $T$ and $\ddot{\phi}$, the density $d(r)$ generated through 5000 Monte Carlo replications. The second row of Figure S.4 (Plots (d), (e), (f)) shows the estimated asymptotic distribution $\widehat{\mathcal{D}}(r)$, namely, the distribution $\mathcal{D}(r)$ where the parameters are estimated from the sample of 5000 Monte Carlo replications. The third row of Figure S.4 (Plots (g), (h), (i)) shows the corresponding $\mathcal{D}(r)$. These were plotted using 5000 values of the argument starting at -1 and increasing by steps of size 0.0004 until 1. As expected, we observe that the degree of approximation of $d(r)$ to $\widehat{\mathcal{D}}(r)$ and to $\mathcal{D}(r)$ improves as $T$ increases and/or $\ddot{\phi}$ decreases. In particular, Plots (a), (d) and (g) in Figure S.4, where $T=50$, show that both $\widehat{\mathcal{D}}(r)$ and $\mathcal{D}(r)$ approximate $d(r)$ well for a low-to-intermediate degree of serial correlation ($\ddot{\phi}\leq 0.36$, i.e. $\phi\leq 0.6$). In contrast, in cases with high degree of serial correlation ($\ddot{\phi}\geq 0.81$, i.e. $\phi\geq 0.9$), $\mathcal{D}(r)$ has larger tails compared to $d(r)$; that is, the latter over-estimates the probability of large spurious correlations. Notably, this is not the case of $\widehat{\mathcal{D}}(r)$, which approximates well $d(r)$ also for $T=50$ and $\ddot{\phi}=0.81$. However, it is noteworthy that the difference between the three densities is negligible for $T\geq 100$ (Figure S.4, Plots (b), (e) and (h) for $T=100$, and Plots (c), (f) and (i) for $T=250$), also with high degree of serial correlation ($\ddot{\phi}\approx 0.90$, i.e. $\phi=0.95$). These numerical experiments corroborate that the sample cross-correlation between orthogonal Gaussian AR(1) processes is affected by the degree of serial correlation in a way that is well approximated by $\mathcal{D}(r)$. In fact, for a sufficiently large finite $T$, we observe that $\Pr\mathopen{}\mathclose{{}\left\{|\widehat{c}_{12}^{x}|\geq\tau}\right\}$, $\tau>0$, increases with $\ddot{\phi}$ in a similar way for $d(r)$, $\widehat{\mathcal{D}}(r)$ and $\mathcal{D}(r)$.
@@
Figure S.4: Monte Carlo densities for $\widehat{c}_{12}^{x}$ (top), corresponding estimated $\widehat{\mathcal{D}}(r)$ (medium) and asymptotic $\mathcal{D}(r)$ (bottom) for various $T$ and $\phi$.
@@
In Section 2.1 of the main text we pointed out that the impact of $\ddot{\phi}$ on $\mathcal{D}(r)$ depends on $Sign(\ddot{\phi})$. In particular, when $-1<\ddot{\phi}<0$, an increment on $|\ddot{\phi}|$ makes the density of $\widehat{c}_{12}^{x}$ more concentrated around 00. In order to validate this result, we run simulations with $T=100$ and different values for the second element of the diagonal of $\mathbf{D}_{\phi}$; namely, $-0.3,-0.6,-0.9,-0.95$. Results are shown in Plots (a) and (b) of Figure S.5. In this case, we see that when $Sign(\phi_{1})\neq Sign(\phi_{2})$ and $|\ddot{\phi}|$ increases, $d(r)$ increases its concentration around 00 in a way that is, again, well approximated by $\mathcal{D}(r)$.
C.2 General Case
To generalize our findings to the case of non-Gaussian weakly correlated AR and ARMA processes, we generate predictors according to the following DGPs: ${x}_{1t}=(\phi+0.1){x}_{1t-1}+(\phi+0.1){x}_{1t-2}-0.2{x}_{1t-3}+u_{1t}$, and ${x}_{2t}=\phi x_{2t-1}+\phi x_{2t-2}+{u}_{2t}+0.8{u}_{2t-1}$, where $t=1,\ldots,100$ and $\phi=0.15,0.3,0.45,0.475$. Moreover, we generate $u_{1t}$ and $u_{2t}$ from a bivariate Laplace distribution with means 00, variances $1$, and $c_{12}^{u}=0.2$. In these more general cases, we do not know an approximate theoretical density for $\widehat{c}_{12}^{u}$. Therefore, we rely entirely on simulations to show the effect of serial correlation on $\Pr\mathopen{}\mathclose{{}\left\{|\widehat{c}_{12}^{x}|\geq\tau}\right\}$. Figure S.6 shows $d(r)$ obtained from 5000 Monte Carlo replications for the different values of $\phi$. In short, also in the more general cases where predictors are non-Gaussian, weakly correlated AR(3) and ARMA(2,1) processes, the probability of getting large sample cross-correlations depends on the degree of serial correlation. More simulation results are provided below.
Figure S.5: Monte Carlo densities for $\widehat{c}_{12}^{x}$ (a) and corresponding $\mathcal{D}(r)$ (b), for $T$=100 and various (negative) $\ddot{\phi}$’s.
@@ C.2 General Case
C.3 More General Cases
@@
The theoretical contribution reported in Section 2.1 requires the Gaussianity of $u_{1}$ and $u_{2}$. With the following simulation experiments we show that the impact of $\ddot{\phi}$ on the density of $\widehat{c}_{12}^{x}$ is relevant also when $u_{1t}$ and $u_{2t}$ are non-Gaussian random variables. To this end, we generate $u_{1t}$ and $u_{2t}$ from the following distributions: Laplace with mean 0 and variance 1 (case (a)); Cauchy with location parameter 0 and scale parameter 1 (case (b)); and from a $t$-student with 1 degree of freedom (case (c)). Figure S.7 reports the results of the simulation experiment. We can state that regardless the distribution of the processes, whenever $Sign(\phi_{1})=Sign(\phi_{2})$, the probability of large values of $\widehat{c}_{12}^{x}$ increases with $\ddot{\phi}$. As a curiosity, this result is more evident for the case of Laplace variables, whereas for Cauchy and $t$-student the effect of $\ddot{\phi}$ declines.
@@
Appendix D More on Comparison with GLS-LASSO
The popular Cochrane-Orcutt GLS estimator allows us to approximate the serial correlation structure of the error term while retaining consistent coefficient estimation (see, e.g.,  Cochrane and Orcutt, 1949). The GLS-LASSO (Chronopoulos et al., 2023) can be summarized in the following steps:
@@ Appendix D More on Comparison with GLS-LASSO
$\widehat{\boldsymbol{\alpha}}=\underset{\boldsymbol{\alpha}\in\mathbb{R}^{n}}{\operatornamewithlimits{argmin\ }}\>\mathopen{}\mathclose{{}\left\{\>\frac{1}{2(T)}\>\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left|\widetilde{\mathbf{y}}-\widetilde{\mathbf{X}}\boldsymbol{\alpha}}\right|}\right|_{2}^{2}+\lambda||\boldsymbol{\alpha}||_{1}\>}\right\},$ (D.12)
@@
The loss function in (D.12) corresponds to the $\ell_{1}$-penalized regression considering the estimates of $\phi_{\varepsilon l}$, $l=1,\dots,p_{\varepsilon}$hronopoulos et al. (2023) provide the theoretical properties of this procedure and support them through simulation results. Thus, the working model of GLS-LASSO is
$y_{t}-\sum_{j=1}^{p_{\varepsilon}}\phi_{\varepsilon j}y_{t-j}=\sum_{i=1}^{n}\alpha_{i}^{*}\mathopen{}\mathclose{{}\left(x_{it}-\sum_{j=1}^{p_{\varepsilon}}\phi_{\varepsilon j}x_{it-j}}\right)+\omega_{t}.$ (D.13)
@@
Common Factor Restriction. The common factor restriction holds when predictors and error term are generated by the same AR($p$) process (Mizon, 1995), as in the Example 1 and Remark 1 of the main text. In this case, we can easily observe that the working model of ARMAr-LASSO (see (5) in the main text) and (D.13) estimate the true coefficients $\alpha_{i}^{*}$ by means of the AR($p$) residuals $u_{it}$. To this end, we consider the simplest case where both predictors and error term are AR(1) processes with autoregression coefficient $\phi$. In this case the GLM-filter leads to $\widetilde{x}_{it}=x_{it}-\phi x_{it-1}=u_{it}$.

### v3

@@ Appendix B Distribution of Sample Correlation Between Serially Correlated Processes: Simulation Experiments
In this Section, we conduct Monte Carlo experiments to assess numerically the approximation of the density of $\widehat{c}_{ij}^{x}$ to $\mathcal{D}(r)$, as described in Section 2.1 of the main text. In particular, we compare the density of $\widehat{c}_{ij}^{x}$ obtained by simulations (indicated as $d(r)$) the distribution provided in Proposition 1 of the main text (indicated as $\mathcal{D}(r)$). After, we expand the theoretical results in more generic contexts, relaxing the assumption that the covariates are orthogonal Gaussian AR(1) processes.
B.1 Numerical Approximation of $d(r)$ to $\mathcal{D}(r)$
@@
We generate data from the bivariate process $\mathbf{x}_{t}=\mathbf{D}_{\phi}\mathbf{x}_{t-1}+\mathbf{u}_{t}$ for $t=1,\dots,T$, where $\mathbf{D}_{\phi}$ is a $2\times 2$ diagonal matrix with the same autocorrelation coefficient $\phi$ in both positions along the diagonal, and $u_{t}\sim N(\boldsymbol{0}_{2},\boldsymbol{I}_{2})$. We consider $T=50,100,250$ and $\phi=0.3,0.6,0.9,0.95$ – thus, the parameter $\ddot{\phi}$ in $\mathcal{D}(r)$, here equal to $\phi^{2}$, takes values 0.09, 0.36, 0.81, 0.90. The first row of Figure S.6 (Plots (a), (b), (c)) shows, for various values of $T$ and $\ddot{\phi}$, the density $d(r)$ generated through 5000 Monte Carlo replications. The second row of Figure S.6 (Plots (d), (e), (f)) shows the corresponding $\mathcal{D}(r)$. These were plotted using 5000 values of the argument starting at -1 and increasing by steps of size 0.0004 until 1. As expected, we observe that the degree of approximation of $d(r)$ to $\mathcal{D}(r)$ improves as $T$ increases and/or $\ddot{\phi}$ decreases. In particular, Plots (a), (d) and (g) in Figure S.6, where $T=50$, show that $\mathcal{D}(r)$ approximates $d(r)$ well for a low-to-intermediate degree of serial correlation ($\ddot{\phi}\leq 0.36$, i.e. $\phi\leq 0.6$). In contrast, in cases with high degree of serial correlation ($\ddot{\phi}\geq 0.81$, i.e. $\phi\geq 0.9$), $\mathcal{D}(r)$ has larger tails compared to $d(r)$; that is, the latter over-estimates the probability of large spurious correlations. However, it is noteworthy that the difference between the two densities is negligible for $T\geq 100$ (Figure S.6, Plots (b), (e) and (h) for $T=100$, and Plots (c), (f) and (i) for $T=250$), also with high degree of serial correlation ($\ddot{\phi}\approx 0.90$, i.e. $\phi=0.95$). These numerical experiments corroborate that the sample cross-correlation between orthogonal Gaussian AR(1) processes is affected by the degree of serial correlation in a way that is well approximated by $\mathcal{D}(r)$. In fact, for a sufficiently large finite $T$, we observe that $\Pr\mathopen{}\mathclose{{\left\{|\widehat{c}_{12}^{x}|\geq\tau}}\right\}$, $\tau>0$, increases with $\ddot{\phi}$ in a similar way for $d(r)$ and $\mathcal{D}(r)$.
@@
Figure S.6: Monte Carlo densities for $\widehat{c}_{12}^{x}$ (top) and asymptotic $\mathcal{D}(r)$ (bottom) for various $T$ and $\phi$.
@@
In Section 2.1 of the main text, we pointed out that the impact of $\ddot{\phi}$ on $\mathcal{D}(r)$ depends on $Sign(\ddot{\phi})$. In particular, when $1<\ddot{\phi}<0$, an increment on $|\ddot{\phi}|$ makes the density of $\widehat{c}_{12}^{x}$ more concentrated around 0. In order to validate this result, we run simulations with $T=100$ and different values for the second element of the diagonal of $\mathbf{D}_{\phi}$; namely, 0.3,0.6,0.9,0.95-0.3,-0.6,-0.9,-0.95. Results are shown in Plots (a) and (b) of Figure S.7. In this case, we see that when $Sign(\phi_{1})\neq Sign(\phi_{2})$ and $|\ddot{\phi}|$ increases, $d(r)$ increases its concentration around 0 in a way that is, again, well approximated by $\mathcal{D}(r)$.
General Case
To generalize our findings to the case of non-Gaussian weakly correlated AR and ARMA processes, we generate covariates according to the following DGPs: ${x}_{1t}=(\phi+0.1){x}_{1t-1}+(\phi+0.1){x}_{1t-2}-0.2{x}_{1t-3}+u_{1t}$, and ${x}_{2t}=\phi x_{2t-1}+\phi x_{2t-2}+{u}_{2t}+0.8{u}_{2t-1}$, where $t=1,\ldots,100$ and $\phi=0.15,0.3,0.45,0.475$. Moreover, we generate $u_{1t}$ and $u_{2t}$ from a bivariate Laplace distribution with means 0, variances $1$, and $c_{12}^{u}=0.2$. In these more general cases, we do not know an approximate theoretical density for $\widehat{c}_{12}^{u}$. Therefore, we rely entirely on simulations to show the effect of serial correlation on $\Pr\mathopen{}\mathclose{{\left\{|\widehat{c}_{12}^{x}|\geq\tau}}\right\}$. Figure S.8 shows $d(r)$ obtained from 5000 Monte Carlo replications for the different values of $\phi$. In short, also in the more general cases where covariates are non-Gaussian, weakly correlated AR(3) and ARMA(2,1) processes, the probability of getting large sample cross-correlations depends on the degree of serial correlation. More simulation results are provided below.
Figure S.7: Monte Carlo densities for $\widehat{c}_{12}^{x}$ (a) and corresponding $\mathcal{D}(r)$ (b), for TT=100 and various (negative) $\ddot{\phi}$’s.
@@
B.2 More General Cases
@@
The theoretical contribution reported in Section 2.1 of the main text requires the Gaussianity of $u_{1}$ and $u_{2}$. With the following simulation experiments we show that the impact of $\ddot{\phi}$ on the density of $\widehat{c}_{12}^{x}$ is relevant also when $u_{1t}$ and $u_{2t}$ are non-Gaussian random variables. To this end, we generate $u_{1t}$ and $u_{2t}$ from the following distributions: Laplace with mean 0 and variance 1 (case (a)); Cauchy with location parameter 0 and scale parameter 1 (case (b)); and from a $t$-student with 1 degree of freedom (case (c)). Figure S.9 reports the results of the simulation experiment. We can state that regardless the distribution of the processes, whenever $Sign(\phi_{1})=Sign(\phi_{2})$, the probability of large values of $\widehat{c}_{12}^{x}$ increases with $\ddot{\phi}$. As a curiosity, this result is more evident for the case of Laplace variables, whereas for Cauchy and $t$-student the effect of $\ddot{\phi}$ declines.
@@
Appendix C Comparison with ARDL and GLS Estimators
The ARDL consists of regressing the response on its past realizations – the autoregressive component – as well as on current and past values of the predictors – the distributed lag component (see, e.g., Panopoulou2004). Although this method does mitigate serial correlation, it has the drawback of requiring a very large number of coefficients to be estimated. This issue becomes particularly relevant when the sample size is small. In contrast, our proposal only requires the addition of a few response lags.
The popular Cochrane-Orcutt GLS estimator approximates the serial correlation structure of the error term while retaining consistent coefficient estimation (see, e.g., CocOrc1949). Although this improves statistical efficiency and inference compared to conventional least squares, it does not tackle directly the risk of spurious correlations due to predictors’ serial correlations, as described in Section 2.1 of the main text. In particular, while the GLS filter may reduce predictors’ serial correlations, it does not remove them completely if the AR structure of the error term differs from the AR or ARMA structures of the predictors. The GLS-LASSO (chronopoulos2023) can be summarized in the following steps:
@@ Appendix C Comparison with ARDL and GLS Estimators
$\widehat{\boldsymbol{\alpha}}=\underset{\boldsymbol{\alpha}\in\mathbb{R}^{n}}{\operatornamewithlimits{argmin\ }}\>\mathopen{}\mathclose{{\left\{\>\frac{1}{2(T)}\>\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\widetilde{\mathbf{y}}-\widetilde{\mathbf{X}}\boldsymbol{\alpha}}}\right|}}\right|_{2}^{2}+\lambda||\boldsymbol{\alpha}||_{1}\>}}\right\},$ (S.15)
@@
The loss function in (S.15) corresponds to the $\ell_{1}$-penalized regression considering the estimates of $\phi_{\varepsilon l}$, $l=1,\dots,p_{\varepsilon}$. chronopoulos2023 provide the theoretical properties of this procedure and support them through simulation results. Thus, the working model of GLS-LASSO is¡
$y_{t}-\sum_{j=1}^{p_{\varepsilon}}\phi_{\varepsilon j}y_{t-j}=\sum_{i=1}^{n}\alpha_{i}^{*}\mathopen{}\mathclose{{\left(x_{it}-\sum_{j=1}^{p_{\varepsilon}}\phi_{\varepsilon j}x_{it-j}}}\right)+\omega_{t}.$ (S.16)
@@
Common Factor Restriction. The common factor restriction holds when predictors and error term are generated by the same AR($p$) process (Mizon1995), as in the Example 1 and Remark 2 of the main text. In this case, we can easily observe that the working model of ARMAr-LASSO (see (1) in the main text) and (S.16) estimate the true coefficients $\alpha_{i}^{*}$ by means of the AR($p$) residuals $u_{it}$. To this end, we consider the simplest case where both predictors and error term are AR(1) processes with autoregression coefficient $\phi$. In this case the GLM-filter leads to $\widetilde{x}_{it}=x_{it}-\phi x_{it-1}=u_{it}$.